# Multiple-Resolution Tokenization for Time Series Forecasting with an Application to Pricing

### v1

@@ 1 Introduction
Forecasting remains a frontier for deep learning models. Non-stationarity, highly dependent stochastic processes, limited predictability, and most importantly small datasets differentiate this prediction problem as a particularly difficult one. The value of improved forecasting capabilities is nearly self-evident as it enables us to improve the contextualisation of decision making. The model we propose was developed specifically for time series for which we have control over some of the auxiliary variables. We assume those auxiliary variables have a significant effect on the time series. Pricing problems are a canonical example in which the auxiliary variable of price affects sales. The ubiquity of pricing problems and resulting large datasets make them a natural candidate for deep learning methods. Markdowns are a subset of pricing problems concerned with reducing prices to clear expiring stock. In this context a forecasting model which incorporates auxiliary variables such as price can be utilised as a simulator to narrow down pricing strategies for empirical testing. There is considerable operational and reputational risk in evaluating previously untrialled strategies. A good simulator enables a virtual exploration of the pricing strategy hypothesis space which can help select potentially more profitable and sustainable pricing strategies for testing.
This work is about tokenisation strategies for time series and aligning a transformer architecture with these strategies. We believe that the main way of advancing transformer capabilities for time series forecasting is by developing and validating specialised tokensiation strategies. Our contributions are:
  • A set of new tokenisation modules to obtain: multiple-resolution past data tokens, multiple-resolution known time-varying tokens, and cross-series tokens. This creates a much broader context window.
  • A reverse splitting output head with favourable scaling properties to deal with the increase in the number of tokens.
  • An application of the model to a real-world forecasting problem faced by the pricing team at a very large retailer. In the experiments we demonstrate that our method outperforms existing in-house methods and two popular existing architectures on a real-world pricing problem.
@@ 2 Deep Learning for Time Series Forecasting
where $t_{0}$ is the point at which we want to make our forecast. The multivariate forecasting problem at $t_{0}$ with a forecast horizon of $f$ is the prediction of future values of the series for the time periods $t_{0}+1$ to $t_{0}+f$ across all channels $c_{1},...,c_{n}\in C$ which we denote as $\hat{y}_{c_{1}:c_{n},t_{0}+1:t_{0}+f}$ to minimise some loss function $loss(y_{c_{1}:c_{n},t_{0}+1:t_{0}+f},\hat{y}_{c_{1}:c_{n},t_{0}+1:t_{0}+f})$. Each channel respectively corresponds to each variate we are interested in forecasting simultaneously. Channels are a choice reflecting our structural view of the time series. Aggregation can be done by concatenating series and adding a static variable denoting some characteristic of the series or by stacking different series along a new dimension.
Time series tend to be non-stationary, that is the joint probability distribution across the same patch resolution is not constant across time. The nature of the data-generating process is fundamentally inconsistent across different time series meaning it is harder to aggregate large corpuses for training. Time series intrinsically produce fewer observations. In most cases we simply do not have a sufficient number of samples for deep learning to be viable. There are limited cases for which we have sufficient corpus sizes in time series: where there is an enormity of related observed processes, a high resolution of measurement, or, a very short forecasting interest relative to the total length of the observed sequence. Given that deep learning requires large datasets applying it to unsuitable forecasting problems is fruitless absent powerful foundation models. The forecasting problem in pricing markdowns is suitable as it benefits from being able to aggregate/concatenate many individual pricing time-series across SKUs and stores.
@@ 2.1 Transformers
Components of the original transformer architecture (Vaswani et al., 2017) powers the state of the art for large language models and has seen more computational resource investment than any other deep learning architecture. Empirically, the transformer is outstanding for language modelling and benefits from scaling laws. Due to the success of transformers with the sequential task of language there has been considerable interest in applying transformers to time series. Academic research has produced mixed results and tends to be limited in the scope of study. In fact there are strong arguments to be made against using transformers for time series with plenty of vocal opponents.
@@
Transformers have architectural limitations when applied to time series (Zeng et al., 2023). Transformers are permutation invariant and thus inherently do not create an inductive bias which reflects that observations happen sequentially (typically in a highly auto-correlated manner). Transformers for language use two core architectural adjustments for dealing with the lack of sequential inductive bias. Positional encoding adds a bias to the latent matrix representation of a language sequence which is designed or learnt to reflect sequentiality. Decoder architectures mask out the attention matrix so that for each unit of latent representation (token) can only observe the tokens that occurred before it in the sequence. The effectiveness of either of these techniques for forecasting is still an open question. The evidence so far is mixed and simple models have been shown to outperform very complex transformer models on academic benchmarks (Zeng et al., 2023).
@@
At a higher level, the concept of a token in language [^1] Note that tokenisation in LLMs is typically not trained end-to-end and is fixed, transferred into the model from a different pretraining task. This is often a cause of problems. does not trivially map to time series. Time series are processes which have regular or irregular observations, but single observations are not necessarily meaningful units in terms of the dynamics of a time series. An analogous reasoning from classical time-series approaches is the concept of the order of integration which is the number of times a time series has to be differenced (how many adjacent points need to be considered) to become stationary. Time series are often decomposed into trend and seasonality or frequencies, both of which are transformations of more than just a single observation. A notable transformer time series architecture uses frequency decomposition to tokenize a time series (Zhou et al., 2022). Alternatively, a popular and increasingly standard method (Das et al., 2023) for tokenisation is patching (Nie et al., 2023): it splits a series up into fixed length sequences which are then mapped into the embedding space using a shared learnable linear transformation. The original PatchTST uses patches of length 16 with a stride of 8 (allowing for overlapping patches). This can be interpreted as the resolution at which a time series is processed. As opposed to previous by observation period encoder-decoder designs (Lim et al., 2021) patching loses the 1-to-1 correspondence between tokens and forecast horizon periods. Instead, the matrix output from the transformer blocks is flattened and passed to a large linear layer (presenting a potential bottleneck) which outputs the forecast vector. The implicit assumption is that meaningful information about the time series occurs at resolutions denser than the base time series. A linear model over the whole series assumes that we need to look at the highest density; the whole series. Our work builds on this assumption by extending patching to multiple resolutions.
@@
2.2 Existing Multiple Resolution Patching Approaches
@@
We would like to outline the differentiating contributions of our work in contrast with other recent approaches which employ the idea of utilising multiple resolution modelling in transformer architectures. We provide a brief description of each model and point out the key difference in capabilities and architecture. Our model is unique in that it uses the tokens from multiple resolutions in the same attention mechanism allowing for explicit pairwise modelling at multiple resolutions across multiple data types. In addition we provide a scheme for how to tokenize time-varying known data at multiple resolutions including for the time steps which occur in the future.
@@
3 The Model: Multiple-Resolution Tokenization (MRT)
Our model is based on two ideas. First, that time series carry relevant information at multiple resolutions and that this should be represented in attention simultaneously. Second, that the resulting multiple resolution tokens can and should be joined with tokens obtained from auxiliary data to form the input matrix for the transformer. This idea is a response to one of the key design questions in transformers for time series forecasting: what are meaningful units of information. Tokens in this context should just be understood as meaningful representations of some facet of the time series. In designing separate tokenisation for past and auxiliary data we implicitly assume that auxiliary data carries significant information. Designs which include auxiliary information often meld all information at a given token together (Lim et al., 2021; Chen et al., 2023), whereas we process separately for each type of auxiliary information. The separation into more tokens enables explicit attention links to aspects of auxiliary information which increases interpretability. The multitude of resolutions complemented by representations of auxiliary variables as a tokenization strategy enables the learning of complex pairwise relationships.
3.1 Multiple-Resolution Patching
Figure 1: Multiple Resolution Patching (MRP)
3.2 Handling Auxiliary Information
Base Tokenization
Time-Varying Known Variables (TVK)
In line with previous multiple resolution patching we apply a modified version built on basis combination. In contrast with past-data we want to patch across tokens (collection of vectors) not individual observations (collection of points). We split the tensor $k_{i}$ roughly equal parts for each resolution $i$ along the time dimension. We use the same resolutions, rules on length (denote the length of a patch $a_{i}$), and padding (so each patch is actually $a_{i}+1$), but for a longer series as the size of $T$ is $l+f$. The splitting leaves us with a number of tensors of the shape $[B,C,L,T=a_{i}+1]$. We treat the columns of each matrix in the last two dimensions of the from $[L=d_{m},T=a_{i}+1]$ as a basis. For each resolution we learn a basis layer: vector of size $a_{i}+1$ which linearly combines the columns of the matrix into a single vector. The basis layer transforms $[B,C,L,T=a_{i}+1]\rightarrow[B,C,L,T=1]=[B,C,L]$ for each patch in each resolution. This operation is illustrated in figure 2. The basis combination results in $\sum_{k_{i}\in K}k_{i}=n_{MRP}$ such tokens. The tensors are stacked along a patch dimension into $[B,C,T=n_{MRP},L]$. To reduce the relative influence of auxiliary variables and reduce the size of the input matrix to the transformer we employ a linear compression layer to compress the patch dimension to a predefined number of tokens $n_{TVK}$. This layer mixes across patch representation. The transformation is a linear layer $[B,C,L,T=n_{MRP}]\rightarrow[B,C,L,T=n_{TVK}]$, which we transpose to $[B,C,T,L]$. This process creates a a set of TVKT which are a mixture of representations extracted at multiple resolutions from joint base representations of the time-varying auxiliary data.
Figure 2: Basis combination operation as used for tokenising time-varying known variables.
Static Variables
3.3 Cross-Series Information
Figure 3: Channel mixer: a module for learning cross-series tokens.
3.4 Output Head
We propose a novel output head architecture which has favourable scaling properties. We call it reverse splitting. The transformer blocks produce a matrix which is the same size as the input matrix $[B,C,T,L]$. Using the same resolution values as for the splitting, we reverse the process. For resolution $k_{i}$ we take the next $k_{i}$ column vectors from the matrix and individually project them into patches that are composed to represent the forecast. The length of the projected patches ($p_{i}$ or $p_{i}+1$) is determined the same way as the splitting in MRP. We learn a projection matrix for each resolution $\mathbb{R}^{d_{m}\times p_{i}+1}$ and omit the final vector value for the reverse patching for which we have set the length to $p$. The core operation iterates over the dimension $T$ so each input is $[B,C,L]$ which is projected into $[B,C,T=p_{i}+1]$, this is repeated $k_{i}$ times and concatenated according to the length rule into $[B,C,T=f]$. This leaves us with $|K|$ forecast vectors of the form $[B,C,T=f]$ which we simply sum to obtain the final output. We call this reverse splitting. We illustrate this module in figure 4. It scales favourably to simple flattening as it requires us to learn $d_{m}p_{i}$ parameters for each resolution $i$ for a total of $d_{m}\sum(p_{i}+1)$. This means that scaling depends on a sum of fractions of the forecast horizon $f$ and inversely depends on the resolution (number of divisions) as $p_{i}\approx f/k_{i}$. Restated, our output head requires approximately $d_{m}f\sum\frac{1}{k_{i}}$ parameters. In comparison, the flattening approach would require learning $d_{m}f\sum k_{i}$ parameters which is considerably greater as $k_{i}$ are natural numbers. Note that the reverse splitting does not use the corresponding position outputs of auxiliary tokens.
Figure 4: Reverse splitter - the output head only operates on post transformer tokens corresponding to the original MRP positions.
3.5 Other Architectural Choices
Masking
Positional Encoding
After Attention
3.6 Architectural Limitations
Multiple resolution patching and associated auxiliary variable tokenisation adds many more hyperparameters in the form of the resolution set. The space of resolution sets is combinatorial making hyperparameter optimisation a challenging task. We somewhat constrain the space by not allowing overlaps, thus avoiding the need for a stride hyperparameter associated with every resolution. Our tokenisation modules increase the number of learnable parameters in the embedding especially for small $k_{i}$ which makes scaling to large latent sizes more difficult. The CST are sequentially computed downstream of all other tokens limiting the degree of parallelisation in tokenisation. Since all tokenisation is done end-to-end a degree of overfitting is expected. Computationally the main difficulty is the increased number of tokens which scales computation quadratically. This constrains how high $n_{MRP}$ can get limiting dense resolution sets which may be useful for longer context windows with long lookback horizons. Exploiting and encouraging sparsity in self-attention may be one possibility to deal with the larger context window.

### v2

@@ 1 Introduction
Forecasting remains a frontier for deep learning models. The value of improved forecasting capabilities is obvious as it enables us to improve the contextualisation of decision making. The model we propose was specifically developed for time series for which we have control over some of the auxiliary variables. We assume those auxiliary variables have a significant effect on the time series. Since auxiliary variables are hard to model with traditional statistical approaches we turned to deep learning. We specifically focus on pricing problems: a canonical example in which the auxiliary variable of price affects sales. The ubiquity of pricing problems and resulting large datasets make them suitable for deep learning models. The forecasting problem in pricing is suitable as it benefits from being able to aggregate many similar individual pricing time-series across stock keeping units (SKUs) and stores. The model we propose is based on the celebrated transformer architecture in which self-attention is leveraged to explicitly model pairwise relationships between tokens. We believe this choice of architecture is justified by the complexity of the underlying process of consumer behaviour in response to changes in price and large dataset sizes.
This work is about tokenisation strategies for time series and aligning a transformer architecture with these strategies. We believe that a key way of advancing transformer capabilities for time series forecasting is by developing and validating specialised tokensiation strategies. Our contributions are:
  • A set of new tokenisation modules to obtain: multiple-resolution (MR) past data tokens, MR known time-varying tokens, and cross-series tokens. We differ from existing multiple-resolution approaches as all tokens are passed to a single attention mechanism creating a larger context window.
  • A reverse splitting output head which encourages output token specialisation and with favourable scaling properties to deal with the increase in the number of tokens.
  • An application of the model to a noisy real-world forecasting problem faced by the pricing team at a very large retailer and ablations on an analogous problem based on a public retail dataset. Our method outperforms existing in-house methods and two popular existing architectures on a both problems.
@@ 2 Deep Learning for Time Series Forecasting
where $t_{0}$ is the point at which we want to make our forecast. The multivariate forecasting problem at $t_{0}$ with a forecast horizon of $f$ is the prediction of future values of the series for the time periods $t_{0}+1$ to $t_{0}+f$ across all channels $c_{1},...,c_{n}\in C$ which we denote as $\hat{y}_{c_{1}:c_{n},t_{0}+1:t_{0}+f}$ to minimise some loss function $loss(y_{c_{1}:c_{n},t_{0}+1:t_{0}+f},\hat{y}_{c_{1}:c_{n},t_{0}+1:t_{0}+f})$. Each channel respectively corresponds to each variate we are interested in forecasting simultaneously. Channels are a choice reflecting our structural view of the time series of what should be predicted simultaneously for a given context. Larger training sets are produced by concatenating series and adding a static variable denoting some characteristic of the series.
@@ 2.1 Transformers
Components of the original transformer architecture (Vaswani et al., 2017) powers the state of the art for large language models and has seen more computational resource investment than any other deep learning architecture. Empirically, the transformer is outstanding for language modelling and benefits from scaling laws. Due to the success of transformers with the sequential task of language there has been considerable interest in applying transformers to time series. Academic research has produced mixed results and tends to be limited in the scope of study.
@@
Transformers have architectural limitations when applied to time series (Zeng et al., 2023). Transformers are permutation invariant and thus inherently do not create an inductive bias which reflects that observations happen sequentially. Transformers for language use two core architectural adjustments for dealing with the lack of sequential inductive bias. Positional encoding adds a bias to the latent matrix representation of a language sequence which is designed or learnt to reflect sequentiality. Decoder architectures mask out the attention matrix so that each unit of latent representation (token) can only observe the tokens that occurred before it in the sequence. The effectiveness of either of these techniques for forecasting is still an open question. The evidence so far is mixed and simple models have been shown to outperform very complex transformer models on academic benchmarks (Zeng et al., 2023).
@@
At a higher level, the concept of a token in language [^1] Note that tokenisation in LLMs is typically not trained end-to-end and is fixed, transferred into the model from a different pretraining task. This is often a cause of problems. does not trivially map to time series. Time series are composed of a sequence of observations over time but single observations are not necessarily meaningful units in terms of the dynamics of a time series. An analogous reasoning from classical time-series approaches is the concept of the order of integration which is the number of times a time series has to be differenced (how many adjacent points need to be considered) to become stationary. Time series are often decomposed into trend and seasonality or frequencies, both of which are transformations of more than just a single observation. A notable transformer time series architecture uses frequency decomposition to tokenize a time series (Zhou et al., 2022). Alternatively, a popular and increasingly standard method (Das et al., 2023) for tokenisation is patching (Nie et al., 2023): it splits a series up into fixed length sequences which are then mapped into a latent space using a shared learnable linear transformation. The original PatchTST uses patches of length 16 with a stride of 8 allowing for overlapping patches. This can be interpreted as the resolution (number of observations informing a token) at which a time series is processed. As opposed to single period to token designs (Lim et al., 2021) patching loses the 1-to-1 correspondence between tokens and forecast horizon periods. Instead, the matrix output from the transformer blocks is flattened and passed to a large linear layer (presenting a potential bottleneck) which outputs the forecast vector. The implicit assumption is that meaningful information about the time series occurs at higher resolutions than the individual observations in the time series. A linear model over the whole series assumes that we need to look at the highest resolution: the whole series. Our work builds on this assumption by extending patching to multiple resolutions.
@@
3 The Model: Multiple-Resolution Tokenization (MRT)
Our model is based on two ideas: that time series carry relevant information at multiple resolutions across all associated data and that all of these relationships should be modelled explicitly. The latter enables us to extract conditional leading indicators which only appear for some subset of series in a large corpus. Transformers are a natural choice for such modelling as they enable us to explicitly model pairwise relationships between tokens. Defining meaningful units of information has been a key question in the development of time series. In this context tokens should be understood as useful representations of some facet of a time series. In designing separate tokenisation for past and auxiliary data we implicitly assume that auxiliary data carries separately significant information. Designs which include auxiliary information often meld all information at a given token together (Lim et al., 2021; Chen et al., 2023), whereas we process separately for each type of auxiliary information. The separation into more tokens enables explicit attention links to aspects of auxiliary information which increases interpretability. The multitude of resolutions complemented by representations of auxiliary variables as a tokenization strategy enables the learning of complex pairwise leading representations.
3.1 Multiple-Resolution Patching
Figure 1: An illustration of Multiple Resolution Patching (MRP). A time series is divided into $k_{i}$ roughly equal parts according to resolutions in $K$. Each part is then mapped to a token.
3.2 Handling Auxiliary Information
Base Tokenization
Time-Varying Known Variables (TVK)
In line with previous multiple resolution patching we apply a modified version built on basis combination. In contrast with past-data we want to patch across tokens (collection of vectors) not individual observations (collection of points). We split the tensor $k_{i}$ roughly equal parts for each resolution $i$ along the time dimension. We use the same resolutions, rules on length (denote the length of a patch $a_{i}$), and padding (so each patch is actually $a_{i}+1$), but for a longer series as the size of $T$ is $l+f$. The splitting leaves us with a number of tensors of the shape $[B,C,L,T=a_{i}+1]$. We treat the columns of each matrix in the last two dimensions of the from $[L=d_{m},T=a_{i}+1]$ as a basis. For each resolution we learn a basis layer: vector of size $a_{i}+1$ which linearly combines the columns of the matrix into a single vector. The basis layer transforms $[B,C,L,T=a_{i}+1]\rightarrow[B,C,L,T=1]=[B,C,L]$ for each patch in each resolution. This operation is illustrated in figure 2. The basis combination results in $\sum_{k_{i}\in K}k_{i}=n_{MRP}$ such tokens. Note that $n_{MRP}$ at this stage is equal for both past data and TVK data despite the two sequences having different lengths as our resolution set is defined in terms of the number of parts a sequence should be split into which corresponds to the number of tokens produced by each resolution. The tensors are stacked along a patch dimension into $[B,C,T=n_{MRP},L]$. To reduce the relative influence of auxiliary variables and reduce the size of the input matrix to the transformer we employ a linear compression layer to compress the patch dimension to a predefined number of tokens $n_{TVK}$. This layer mixes across patch representation. The transformation is a linear layer $[B,C,L,T=n_{MRP}]\rightarrow[B,C,L,T=n_{TVK}]$, which we transpose to $[B,C,T,L]$. This process creates a a set of TVKT which are a mixture of representations extracted at multiple resolutions from joint base representations of the time-varying auxiliary data.
Figure 2: Basis combination operation as used for tokenising time-varying known variables. A matrix is divided along the time dimension according to the resolution set $K$. The columns in each patch are combined using a resolution-wise learnt linear combination rule.
Static Variables
3.3 Cross-Series Information
Figure 3: Channel mixer: a module for learning cross-series tokens. The channel mixer takes all other tokens and passes them through a mixer architecture which applies linear layers across the time and then the channel dimension.
3.4 Output Head
We propose reverse splitting, a novel output head which has favourable scaling properties. The transformer blocks produce a matrix which is the same size as the input matrix $[B,C,T,L]$. Using the same resolution values as for the splitting, we reverse the process. For resolution $k_{i}$ we take the next $k_{i}$ column vectors from the output matrix corresponding to the input positions of the MRP tokens and individually project them into patches that are then composed to represent the forecast. The length of the projected patches ($p_{i}$ or $p_{i}+1$) is determined the same way as the splitting in MRP. We learn a projection matrix for each resolution $\mathbb{R}^{d_{m}\times p_{i}+1}$ and omit the final vector value for the reverse patching for which we have set the length to $p$. The core operation iterates over the dimension $T$ so each input is $[B,C,L]$ which is projected into $[B,C,T=p_{i}+1]$, this is repeated $k_{i}$ times and concatenated according to the length rule into $[B,C,T=f]$. This leaves us with $|K|$ forecast vectors of the form $[B,C,T=f]$ which we simply sum to obtain the final output. We call this reverse splitting. We illustrate this module in figure 4. It scales favourably to simple flattening as it requires us to learn $d_{m}p_{i}$ parameters for each resolution $i$ for a total of $d_{m}\sum(p_{i}+1)$. This means that scaling depends on a sum of fractions of the forecast horizon $f$ and inversely depends on the resolution (number of divisions) as $p_{i}\approx f/k_{i}$. Restated, our output head requires approximately $d_{m}f\sum\frac{1}{k_{i}}$ parameters. In comparison, the flattening approach would require learning $d_{m}f\sum k_{i}$ parameters which is considerably greater as $k_{i}$ are natural numbers. Note that the reverse splitting does not use the corresponding position outputs of auxiliary tokens and can be applied even if auxiliary information is not included. In addition to favourable scaling, the design of reverse splitting was motivated by enabling the learning of forecast mappings at different resolutions. These resolutions have a symmetry with MRP patching and enable the model to learn a multiplicity of output mappings increasing predictive robustness similarly to ensembles. Moreover, output tokens are specialised to predict a segment of a specific length at a specific position so each output token explicitly serves a different purpose. This is in contrast to typical output head designs in which the output tokens are flattened and passed to a single linear layer meaning there is representational multiplicity.
Figure 4: An illustration of the reverse splitter output head. The reverse splitter only operates on post transformer tokens corresponding to the original MRP positions. Each token is projected into a proportional part of the final forecast according to the corresponding resolution from the resolution set $K$. The forecasts produced by the different resolutions are then summed together to produce the final forecast.
3.5 Other Architectural Choices
Masking
Positional Encoding
After Attention
3.6 Limitations
Multiple resolution patching and associated auxiliary variable tokenisation adds many more hyperparameters in the form of the resolution set. The space of resolution sets is combinatorial making hyperparameter optimisation a challenging task. We somewhat constrain the space by not allowing overlaps, thus avoiding the need for a stride hyperparameter associated with every resolution. Our tokenisation modules increase the number of learnable parameters in the embedding especially for small $k_{i}$ which makes scaling to large latent sizes more difficult. The CST are sequentially computed downstream of all other tokens limiting the degree of parallelisation in tokenisation. All tokenisation is done end-to-end a degree of overfitting is expected especially with CST tokens as they are computed using all other tokens.
Computationally the main difficulty is the increased number of tokens which scales computation quadratically. This constrains how high $n_{MRP}$ can get limiting dense resolution sets (high $k_{i}$) which we hypothesise are useful as longer context windows for problems with long lookback horizons. Our novel output head design provides a counterbalance in terms of scaling as it is much more computationally efficient than flattening output heads which tend to consume the bulk of computation in models like PatchTST. Exploiting and encouraging sparsity in self-attention may be one possibility to deal with the larger context window. Current production large language models are designed to handle context windows which are three to five orders of magnitude higher than what we demonstrate in our experiments.
4 Existing Multiple Resolution Approaches
@@
We would like to outline the differentiating contributions of our work in contrast with other recent approaches which employ the idea of utilising multiple resolution modelling in transformer architectures. We provide a brief description of each model and point out the key difference in capabilities and architecture. Our model is unique in that it uses the tokens from multiple resolutions in the same attention mechanism allowing for explicit pairwise modelling at multiple resolutions across multiple data types and uses a completely novel output which encourages every relevant output token to assume a specialised role. In addition we provide a scheme for how to tokenize time-varying known data at multiple resolutions including for the time steps which occur in the future.
@@
Note on Other Multi-Scale Approaches
There is a host of other modelling approaches which use a multiplicity of some input to extract and process representations. Multi-fidelity modelling (Perdikaris et al., 2017) has been used to infer relationships between cheap to obtain low fidelity data and expensive to obtain high fidelity data. The main interpolation engine used are Gaussian processes which are well suited as universal approximators but have cubic complexity in the number of samples rendering them impractical for large scale problems. They are used in artificially enhancing fidelity in physical simulation. Neural operators (Kovachki et al., 2023) learn maps between continuous infinite-dimensional functional spaces taking the idea of multi-fidelity modelling to the limit as they can take any discretization/level of fidelity as an input and can be queried at any level of discretization for an output. A canonical application is learning surrogate maps for the solution operators of PDEs and weather forecasting Kurth et al. (2023). The capability is impressive but adds an unnecessary level of complexity as output discretization is fixed in our problem. In terms of input fidelity levels, we artificially induce multiple levels of fidelity through tokenisation and have control over this as a hyperparameter so being fully flexible at the input level is not necessary. The fixed structure of the forecasting problem of regularly spaced observations at a single frequency means the additional capabilities of multi-fidelity modelling are not necessary and would add an undesirable level of complexity.
Mixed-Resolution approaches partitions the input into a single set where the partitions can be of different sizes. The partitioning in vision applications (Ronen et al., 2023) is determined by a saliency algorithm. So long as resolution sizes match all partitions present in a mixed-resolution approach will also appear in MRT. Mixed-resolution can be interpreted as a masked MRT with specific constraints on the masking such as the unmasked partitions composing exactly one input according to a saliency score. While this preselection is interesting, it assumes that we should only extract information at a single resolution for each partition of the input limiting the representations we can extract as relevant dynamics may exist between pairs of less salient frequencies.